### v1

@@ Abstract
Image segmentation plays a crucial role in extracting important objects of interest from images, enabling various applications. While existing methods have shown success in segmenting clean images, they often struggle to produce accurate segmentation results when dealing with degraded images, such as those containing noise or occlusions. To address this challenge, interactive segmentation has emerged as a promising approach, allowing users to provide meaningful input to guide the segmentation process. However, an important problem in interactive segmentation lies in determining how to incorporate minimal yet meaningful user guidance into the segmentation model. In this paper, we propose the quasi-conformal interactive segmentation (QIS) model, which incorporates user input in the form of positive and negative clicks. Users mark a few pixels belonging to the object region as positive clicks, indicating that the segmentation model should include a region around these clicks. Conversely, negative clicks are provided on pixels belonging to the background, instructing the model to exclude the region near these clicks from the segmentation mask. Additionally, the segmentation mask is obtained by deforming a template mask with the same topology as the object of interest using an orientation-preserving quasiconformal mapping. This approach helps to avoid topological errors in the segmentation results. We provide a thorough analysis of the proposed model, including theoretical support for the ability of QIS to include or exclude regions of interest or disinterest based on the user’s indication. To evaluate the performance of QIS, we conduct experiments on synthesized images, medical images, natural images and noisy natural images. The results demonstrate the efficacy of our proposed method.
@@ 1 Introduction
Image segmentation is a fundamental task in computer vision that involves extracting the objects of interest from an image. This process has significant applications in various fields, including medical imaging, autonomous driving, and image understanding. Nevertheless, image segmentation becomes especially challenging when images are degraded due to factors such as noise, blurry boundaries, or occlusions. To address these challenges, prior-based segmentation methods have emerged as effective solutions. Prior-based segmentation incorporates prior information, such as geometric or topological constraints, into the segmentation process. Among the diverse range of prior-based segmentation methods, interactive segmentation stands out as a particularly effective approach. Interactive segmentation leverages the collaborative strengths of humans and machines, allowing users to provide interactive input to guide the segmentation process. Through techniques such as mouse clicks or scribbles, users actively participate in refining the segmentation and providing valuable feedback. This fusion of human expertise and mathematical models has demonstrated significant enhancements in the precision and effectiveness of image segmentation tasks.
@@
Interactive segmentation possesses several distinct advantages. It permits the adjustment of segmented masks based on user-specific requirements. This flexibility expands the applicability of many segmentation models. Moreover, since interactive segmentation can incorporate human guidance, its segmentation results become more versatile and believable in various scenarios. For example, in clinical medicine, delineating diseased regions accurately is a common requirement. However, due to differences in imaging device configurations, direct outputs from a non-interactive model may not yield satisfactory results for images captured with various devices. Improvements and corrections are often necessary to achieve superior segmentation. Instead of overhauling algorithms to suit the unique setup of each imaging device, an interactive segmentation model, capable of accommodating specific user guidance, offers an economical and readily deployable solution. This approach instills greater confidence in clinical professionals as they can actively participate in the segmentation process.
However, while interactive segmentation has attracted extensive attention from various research groups, it still faces challenges and limitations that hinder its widespread adoption. Learning-based methods, such as Segment Anything, require a substantial amount of training data to effectively label images. This reliance on data availability may limit their applicability when only a limited amount of ground truth data is accessible. Additionally, incorporating user inputs into optimization models remains a significant challenge for optimization-based approaches. Many of these methods lack a strong theoretical foundation, which can significantly impact their effectiveness in accurately incorporating user guidance. For instance, the parameters used in the term that integrates user inputs often lack theoretical justification. The selection and adjustment of these parameters, such as weighting factors, are often based on empirical heuristics rather than well-established theoretical principles. This lack of theoretical grounding can result in suboptimal integration of user input and may limit the overall performance and reliability of interactive segmentation methods. Furthermore, existing interactive segmentation approaches often result in outliers and incorrect topology, necessitating additional operations to correct these errors. Therefore, further advancements are needed to develop an accurate and robust theoretical framework to guide the incorporation of user guidance for advancing the effectiveness and accuracy of interactive segmentation.
@@
In this paper, we present our method, called the Quasi-Conformal Interactive Segmentation (QIS), as a solution to the aforementioned challenges. QIS is a deformable model designed for image segmentation, utilizing the fidelity term of the Mumford-Shah model [35]. It achieves segmentation by optimizing a quasi-conformal mapping that can effectively warp a template mask to fit a desired region of interest. The core of our approach relies on a quasi-conformal registration-based segmentation, which forms the backbone of our method. User interactions are defined by two actions: a positive click and a negative click, represented by left and right clicks, respectively. These clicks serve to indicate the regions that the user wants to include or exclude from the segmentation process. To incorporate user input seamlessly, we transform the clicks into a click map, which is then integrated into the proposed model using a novel interactive segmentation energy term that we have designed. The interactive segmentation energy function has been carefully analyzed and designed to ensure that the model accurately includes or excludes the regions of interest or disinterest indicated by the click map associated with each click. Detailed explanations of our method and the energy term can be found in Section 5.
@@ 2.1 Computational quasi-conformal geometry
Computational quasi-conformal geometry serves as a mathematical tool for analyzing and managing geometric distortions inherent in mappings. Notably, conformal mappings, a subset of quasi-conformal mappings, have proven invaluable in geometry processing and have been extensively utilized in tasks such as texture mapping and surface parameterizations [18, 19, 30]. The quantitative assessment of local geometric distortions in a mapping often involves the application of the associated Beltrami coefficient. By manipulating these coefficients, precise control over the geometric characteristics of the mapping can be achieved. This manipulation has led to the development of surface parameterization methods that minimize conformality distortion, with the Beltrami coefficient playing a central role in their formulation [8, 9].
Beyond surface parameterizations, quasi-conformal mappings find applications in computational fabrication [13, 36, 46]. The versatility of quasi-conformal geometry is further underscored by the introduction of various quasi-conformal imaging models in recent years, addressing diverse imaging tasks. These models have been applied in image registration [26, 34], surface matching [10], and shape-prior image segmentation [43, 52]. The utilization of computational quasi-conformal geometry thus stands as a key element in the advancement of methodologies across a spectrum of imaging-related domains.
@@ 2.2 Interactive segmentation
Before the era of deep learning, researchers [17, 20, 25, 42] take interactive segmentation as an optimization or clustering procedure. Badshah and Chen [38] proposed a selective segmentation method using active contour approach. Then, a variational selective segmentation model named as Rada-Chen model is formulated and solved using level-set approach [37, 38]. Spencer and Chen [47] proposed a new model where a global minimizer can be found independently of initial user input, and is further improved by a multigrid algorithm based on a hybrid smoother [41]. To reduce the sensitivity on the user input in a selective segmentation model, Spencer et al. [48] proposed a parameter-free model. Based on a variational segmentation model, Roberts et al. [40] utilized the edge-weighted geodesic distance from a marker set as a penalty term. Ali et al. [1] incorporates an area-based fitting term into a variational selective segmentation model for multi-region segmentation.
DIOS [49] first introduces deep learning into interactive segmentation by embedding positive and negative clicks into distance maps, and concatenating them with the original image as input. It formulates the primary pipeline and train/val protocol for click-based interactive segmentation. After this, [31, 32] focus on the issue of ambiguity and predict multiple potential results and let a selection network or the user choose from them. FCANet [33] emphasizes the particularity of the first click and uses it to construct visual attention. BRS [23] first introduces online optimization, which enables the model to update during annotation. f-BRS [44] speeds up the BRS [23] by executing online optimization in some specific layers. CDNet [6] introduces self-attention into interactive segmentation to predict more consistent results. RITM [45] adds the previous mask as network input to make the prediction more robust and accurate. FocalClick [7] first predicts a coarse segmentation on the Target Crop, and then refines locally on its Focus Crop. These methods achieve satisfactory performances but lack of theoretical analysis of the segmentation result.
@@ 3.1 Mathematical Background on Quasi-Conformal Geometry
$\frac{\partial f}{\partial\bar{z}}=\mu(z)\frac{\partial f}{\partial z}$ (1)
@@
$\begin{split}|\mu(z)|^{2}&=\left|\frac{\partial f}{\partial\bar{z}}/\frac{\partial f}{\partial z}\right|^{2}\\ &=\frac{(\frac{\partial f_{1}}{\partial x_{1}}-\frac{\partial f_{2}}{\partial x_{2}})^{2}+(\frac{\partial f_{2}}{\partial x_{1}}+\frac{\partial f_{1}}{\partial x_{2}})^{2}}{(\frac{\partial f_{1}}{\partial x_{1}}+\frac{\partial f_{2}}{\partial x_{2}})^{2}+(\frac{\partial f_{2}}{\partial x_{1}}-\frac{\partial f_{1}}{\partial x_{2}})^{2}}\\ &=\frac{\|\nabla f\|_{\mathrm{F}}^{2}-2\mathrm{det}\nabla f}{\|\nabla f\|_{\mathrm{F}}^{2}+2\mathrm{det}\nabla f},\end{split}$ (2)
@@
In addition, there is a one-to-one correspondence between the Beltrami coefficient $\mu$ and the quasi-conformal mapping $f$. On one hand, by (1), the Beltrami coefficient $\mu$ can be uniquely calculated if a quasi-conformal mapping $f$ is given. On the other hand, if a Beltrami coefficient $\mu$ is given, the existence and uniqueness of the corresponding quasi-conformal mapping $f$ is established in the following theorem.
Theorem 3.1 (measurable Riemann mapping theorem [16]).
Suppose $\mu:\mathbb{C}\rightarrow\mathbb{C}$ is Lebesgue measurable satisfying $\|\mu\|_{\infty}<1$, then there exists a quasi-conformal mapping $f:\mathbb{C}\rightarrow\mathbb{C}$ in the Sobolev space $W^{1,2}$ that satisfies the Beltrami equation (1) in the distribution sense. Further, assuming that the mapping is stationary at $0,1$, and $\infty$, then the associated quasi-conformal mapping $f$ is uniquely determined.
Further, denote $\sigma_{1}\geq\sigma_{2}>0$ as the singular values of $\nabla f$. Then by (2) again, we have $|\mu(z)|=\frac{\sigma_{1}-\sigma_{2}}{\sigma_{1}+\sigma_{2}}$. Next, we define by $K(z)$ the dilatation
$K(z)=\frac{1+|\mu(z)|}{1-|\mu(z)|}$ (4)
@@ 3.2 Quasi-conformal registration-based segmentation model
The registration-based segmentation model is to extract the interested region by employing the registration framework. For doing the registration, we require two images, called the template and reference. We can take the target image as the template and prescribe a prior image as the reference, where the structure of the prior image is related to the target image. For example, if the target image has only one object, we can set a disk as the prior image. Ideally, after doing the registration, the boundary of the object in the target image can be linked to the boundary of the object in the prior image by the resulting transformation. Since the prior image is given by users and the location of the boundary of the object in the prior image is known, we can identify the boundary of the interested region in the target image. In addition, if the resulting transformation is one-to-one, a topology-preserving segmentation is guaranteed, which means that the topological structure of the segmentation result is the same with the topological structure of the prior image.
Denote $I(\bm{x}):\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}$ as the target image. For simplicity, we consider two-phase segmentation, namely, segmenting the image domain into the foreground and background. In this way, we can define the prior image as
@@
$\mathcal{X}_{D}(\bm{x})=\begin{cases}1,\quad\bm{x}\in D\\ 0,\quad\mathrm{otherwise}\end{cases}$
@@
$\min_{c_{1},c_{2},\varphi(\bm{x})}\frac{1}{2}\int_{\Omega}(I(\varphi(\bm{x}))-J(c_{1},c_{2},\bm{x}))^{2}\mathrm{d}\bm{x}+\alpha\mathcal{R}(\varphi(\bm{x})),$
@@
$\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I-J(c_{1},c_{2},\varphi))^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x},$ (5)
@@
where $\Delta$ is the Laplace operator, $|\mu(\varphi)|^{2}$ is computed by (2), $\phi(v)$ is a scaling function defined for positive $v$ which will be chosen in Section 6. This scaling function should scale up the value of $v$ when it approachs $1$ and scale down if it’s near zero. Through this scaling, the value of $v$ should be regularized to be smaller than $1$ as the energy will be dramatically added when $v$ goes toward $1$. $\alpha_{1}$ and $\alpha_{2}$ are nonnegative parameters weighting the two regularization terms. We can see that the second term is to control the smoothness and the third term is to restrict the distortion.
We also want to emphasize that we can interchange the position of the template and reference and obtain an equivalent model
@@
$\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I(\varphi)-J(c_{1},c_{2}))^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x}.$ (6)
Here, the model (6) is easier to be solved than the model (5) because $c_{1}$, $c_{2}$ and $\varphi$ are decoupled. However, the analysis of the model (5) is more straightforward, which will be illustrated in the following subsection.
4 Revisit on The Fidelity Term of (5)

### v2

@@ Abstract
Interactive segmentation allows users to provide meaningful input to guide the segmentation process. However, an important problem in interactive segmentation lies in determining how to incorporate minimal yet meaningful user guidance into the segmentation model. In this paper, we propose the quasi-conformal interactive segmentation (QIS) model, which incorporates user input in the form of positive and negative clicks. Users mark a few pixels belonging to the object region as positive clicks, indicating that the segmentation model should include a region around these clicks. Conversely, negative clicks are provided on pixels belonging to the background, instructing the model to exclude the region near these clicks from the segmentation mask. By solving our proposed theoretical supported model, the segmentation mask is obtained by deforming a template mask with the same topology as the object of interest using a quasiconformal mapping. This approach makes each user input effectively used and helps to avoid topological errors in the segmentation results. We provide a thorough theoretical analysis of the proposed model for its ability to include or exclude regions of interest or disinterest based on the user’s indication. To evaluate the performance of QIS, we conduct experiments on synthesized images, medical images, and natural images. The results demonstrate the efficacy of our proposed method.
@@ 1 Introduction
Interactive segmentation leverages the collaborative strengths of humans and machines, allowing users to provide interactive input to guide the segmentation process[38, 34]. Through techniques such as mouse clicks or scribbles, users actively participate in refining the segmentation and providing valuable feedback. As a human-involved technique, the segmented results of interactive segmentation become more versatile and believable in various scenarios. For example, in clinical medicine, accurately delineating diseased regions is a common requirement. However, due to differences in imaging device configurations, direct outputs from a non-interactive model may not yield satisfactory results for images captured with various devices. Improvements and corrections are often necessary to achieve superior segmentation. Instead of overhauling algorithms to suit the unique setup of each imaging device, an interactive segmentation model, capable of accommodating specific user guidance, offers an economical and readily deployable solution. This approach instills greater confidence in clinical professionals as they can actively participate in the segmentation process.
An efficient interactive segmentation model should minimize the need for user guidance. As such, three key principles should be followed for designing an effective interactive segmentation model: (1)Effective interaction response: Each new interaction input should result in a meaningful segmentation correction. For example, if a user clicks to include a specific region, the model should at least encompass the local area around the click. Otherwise, the click becomes ineffective, making the correction process cumbersome and unproductive. Despite its intuitive nature, many existing methods fail to achieve this, as highlighted in our experimental section. (2)Adaptability: A robust interactive segmentation model should be easily adaptable to various image modalities and types. In contrast, learning-based methods often require extensive training data with well-labeled ground truth, limiting their applicability to new image types. Given that labeling images is a primary and common application scenario for interactive segmentation, the reliance on high-quality training data is a significant drawback. (3)Minimal outliers: A good interactive segmentation model should minimize the production of noise and outliers. If the model generates a substantial amount of noise, it becomes tedious and time-consuming to manually remove these inaccuracies, ultimately defeating the purpose of computer-assisted segmentation by bringing extra workloads.
@@
Following those principles, in this paper, we present our method, called the Quasi-Conformal Interactive Segmentation (QIS), as a solution to the aforementioned challenges. QIS is a deformable model designed for image segmentation, utilizing the fidelity term of the Mumford-Shah model [32]. It achieves segmentation by optimizing a quasi-conformal mapping that can effectively warp a template mask to fit a desired region of interest. The core of our approach relies on a quasi-conformal registration-based segmentation, which forms the backbone of our method. User interactions are defined by two actions: a positive click and a negative click, represented by left and right clicks, respectively. These clicks serve to indicate the regions that the user wants to include or exclude from the segmentation process. To incorporate user input seamlessly, we transform the clicks into a click map, which is then integrated into the proposed model using a novel interactive segmentation energy term that we have designed. The interactive segmentation energy function has been carefully analyzed and designed to ensure that the model accurately includes or excludes the regions of interest or disinterest indicated by the click map associated with each click. Detailed explanations of our method and the energy term can be found in Section 4.
@@ 2.1 Computational quasi-conformal geometry
Computational quasi-conformal geometry is a mathematical tool for analyzing and managing geometric distortions in mappings. Conformal mappings, a subset of quasi-conformal mappings, are crucial in geometry processing, particularly in tasks like texture mapping and surface parameterizations [16, 17, 27, 31]. The Beltrami coefficient quantitatively assesses local geometric distortions and allows precise control over mapping characteristics, aiding in minimizing conformality distortion [8, 9].
@@ 2.2 Interactive segmentation
Interactive segmentation is typically formulated as an optimization problem [15, 18, 23]. Rother et al. developed GrabCut by solving an optimization problem over graph representation [38]. Badshah and Chen [34] developed a selective segmentation method using active contours. [33, 34] proposed the Rada-Chen model, solved with a level-set approach. Spencer and Chen [42, 43] introduced a model with a global minimizer independent of initial user input, later improved by a multigrid algorithm [37]. Roberts et al. [36] incorporated edge-weighted geodesic distance, while Ali et al. [1] added an area-based fitting term for multi-region segmentation. DIOS [45] pioneered deep learning in interactive segmentation, embedding clicks into distance maps and using them with the original image. Subsequent works [28, 29] addressed ambiguity by predicting multiple potential results. FCANet [30] used the first click for visual attention construction. BRS [21] introduced online optimization for model updates during annotation, which f-BRS [40] accelerated by targeting specific layers. CDNet [7] employed self-attention for consistency, and RITM [41] added the previous mask as input for robustness.
@@ 3.1 Mathematical background on quasi-conformal geometry
$\frac{\partial f}{\partial\bar{z}}=\mu(z)\frac{\partial f}{\partial z}$
@@
(1) $\begin{split}|\mu(z)|^{2}&=\left|\frac{\partial f}{\partial\bar{z}}/\frac{\partial f}{\partial z}\right|^{2}\\ &=\frac{(\frac{\partial f_{1}}{\partial x_{1}}-\frac{\partial f_{2}}{\partial x_{2}})^{2}+(\frac{\partial f_{2}}{\partial x_{1}}+\frac{\partial f_{1}}{\partial x_{2}})^{2}}{(\frac{\partial f_{1}}{\partial x_{1}}+\frac{\partial f_{2}}{\partial x_{2}})^{2}+(\frac{\partial f_{2}}{\partial x_{1}}-\frac{\partial f_{1}}{\partial x_{2}})^{2}}\\ &=\frac{\|\nabla f\|_{\mathrm{F}}^{2}-2\mathrm{det}\nabla f}{\|\nabla f\|_{\mathrm{F}}^{2}+2\mathrm{det}\nabla f},\end{split}$
@@
Further, denote $\sigma_{1}\geq\sigma_{2}>0$ as the singular values of $\nabla f$. Then by (1) again, we have $|\mu(z)|=\frac{\sigma_{1}-\sigma_{2}}{\sigma_{1}+\sigma_{2}}$. We define by $K(z)$ the dilatation
$K(z)=\frac{1+|\mu(z)|}{1-|\mu(z)|}$
@@ 3.2 Quasi-conformal registration-based segmentation model
The registration-based segmentation model is to extract the interested region by employing the registration framework. For doing the registration, we require template and reference. We can prescribe a prior image as the template and take the target image as the reference, where the structure of the prior image is related to the target image. Ideally, after doing the registration, the boundary of the object in the target image can be linked to the boundary of the object in the prior image by the resulting transformation. In addition, if the resulting transformation is one-to-one, a topology-preserving segmentation is guaranteed, which means that the topological structure of the segmentation result is the same with the topological structure of the prior image.
Denote $I(\bm{x}):\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}$ as the target image. For simplicity, we consider a two-phase segmentation, namely, segmenting the image domain into the foreground and background. In this way, we can define the prior image as
@@
$\mathcal{X}_{D}(\bm{x})=\begin{cases}1,\quad\bm{x}\in D,\\ 0,\quad\mathrm{otherwise},\end{cases}$
@@
$\min_{c_{1},c_{2},\varphi(\bm{x})}\frac{1}{2}\int_{\Omega}(I(\bm{x})-J(c_{1},c_{2},\varphi(\bm{x})))^{2}\mathrm{d}\bm{x}+\alpha\mathcal{R}(\varphi(\bm{x})),$
@@
(3) $\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I-J(c_{1},c_{2},\varphi))^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x},$
@@
where $\Delta$ is the Laplace operator, $|\mu(\varphi)|^{2}$ is computed by (1), and $\phi(v)$ is a scaling function defined for positive $v$ which will be chosen in Section 5. This scaling function should scale up the value of $v$ when it approaches $1$ and scale down if it is near zero. Through this scaling, the value of $v$ should be regularized to be smaller than $1$ as the energy will be dramatically increased when $v$ goes toward $1$. $\alpha_{1}$ and $\alpha_{2}$ are nonnegative parameters weighting the two regularization terms. We can see that the second term is to control the smoothness and the third term is to restrict the distortion.
4 Method
In this section, we detail our proposed Quasi-conformal Interactive Segmentation (QIS) model. As an interactive model, our model solves a quasi-conformal registration-based segmentation model (3) in each step. By employing the quasi-conformal approach, the mapping for the deformable segmentation model is guaranteed to be bijective, thus preserving the topology. This control over topology helps avoid noise and outliers in the results. Additionally, we provide theoretical analyses that support our model’s ability to efficiently incorporate user input. These analyses demonstrate that with each new click, the local region around the click will be accurately included or excluded, ensuring an effective and precise segmentation.
4.1 Overall model
Figure 2: A flow chart illustrating the process of the quasi-conformal interactive segmentation (QIS) model.
In this subsection, we first provide an overview of the interactive segmentation procedure. Figure 2 visually illustrates the process. Initially, the quasi-conformal registration-based segmentation model, described in (3), is applied to obtain an initial segmentation. This initial segmentation serves as the current segmentation mask, which may not be entirely accurate and thus requires further adjustments. To guide these adjustments, the user can indicate regions to be included (using positive clicks) and regions to be excluded (using negative clicks). These clicks are then converted into a click map $M_{\mathrm{clicks}}$, which is subsequently incorporated into the quasi-conformal segmentation model. By minimizing the interactive segmentation energy, a new quasi-conformal map is generated, resulting in an improved segmentation mask. If the improved segmentation is still not satisfactory, the process is repeated until a satisfactory segmentation result is achieved. This iterative procedure, which incorporates the user’s input, is referred to as the interactive segmentation process.
Interactive segmentation involves iteratively updating the segmented region until it encompasses the entire region of interest. As a result, the overall model can be divided into multiple steps, with each step focusing on introducing an appropriate click map $M_{\mathrm{clicks}}$ and segmenting a new image. Therefore, the overall process for our interactive segmentation model can be outlined as follows.
Step 0. Given a initial mask $D$ and parameters $\alpha_{1}$ and $\alpha_{2}$, set $I^{0}=I$ and compute $\varphi^{0}$ by solving the following model:
@@ 4.1 Overall model
$\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I^{0}-c_{1}\mathcal{X}_{D}\circ\varphi-c_{2}\mathcal{X}_{D^{c}}\circ\varphi)^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x}.$
Step 1. Give the positive/negative click, compute the click map $M^{1}_{\mathrm{clicks}}$ by (5), choose a suitable weight $r^{1}$ by (16) or (17) and set $I^{1}=I^{0}+r^{1}M^{1}_{\mathrm{clicks}}$. Then given $\varphi^{0}$ as the initial guess, compute $\varphi^{1}$ by solving the following model:
$\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I^{1}-c_{1}\mathcal{X}_{D}\circ\varphi-c_{2}\mathcal{X}_{D^{c}}\circ\varphi)^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x}.$
$\vdots$
Step n. Give the positive/negative click, compute the click map $M^{n}_{\mathrm{clicks}}$ by (5), choose a suitable weight $r^{n}$ by (16) or (17) and set $I^{n}=I^{n-1}+r^{n}M^{n}_{\mathrm{clicks}}$. Then given $\varphi^{n-1}$ as the initial guess, compute $\varphi^{n}$ by solving the following model:
(4) $\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I^{n}-c_{1}\mathcal{X}_{D}\circ\varphi-c_{2}\mathcal{X}_{D^{c}}\circ\varphi)^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x}.$
Note that in each step of the overall model, we solve a quasi-conformal registration-based segmentation model (3). This model can be regarded as the backbone of the whole method. The topology-preserving of the backbone provides great efficiency and helps to segment the images affected by noise or degraded by overexposure or underexposure [48, 49]. Such advantages brought by the topology-preserving property are further illustrated by the improvement of the segmentation accuracy given in Section 6. The segmentation results are iteratively refined based on user input, with each click transforming into a click map. This click map, along with a carefully selected parameter $r$ (discussed in Section 4.3.2), is used to update $I^{n-1}$ to $I^{n}$ as outlined earlier. The segmentation mask at step $n$ is then derived by solving the sub-problem specified in equation (4). This iterative process continues until the segmentation results are deemed satisfactory.
In the following subsections, we will discuss the definitions of the positive click, negative click, and click the map in detail. Additionally, we will provide a comprehensive explanation of the interactive segmentation fidelity term, which will reveal how to choose the weight $r$.
4.2 Click map
In this subsection, we introduce the positive click, negative click, and click map, which is essential for constructing our proposed interactive segmentation model.
Figure 3: An example demonstrating the division of isolated regions in $\mathcal{K}_{i}$ into multiple $D^{i}_{j}$ for $k=3$. In this example, the image domain is classified into three distinct regions based on pixel values: (1) background white, (2) blue, and (3) yellow. Subsequently, the disconnected regions within each class $\mathcal{K}_{i}$ are further divided into individual $D^{i}_{j}$’s.
In this work, we propose a methodology to generate the click map from user clicks based on $K$-means clustering. Suppose a collection $\mathcal{C}_{*}$ of positive clicks or negative clicks is given. In other words, $\mathcal{C}_{*}=\mathcal{C}_{p}$ or $\mathcal{C}_{n}$. Illustrated by Figure 3. the image domain $\Omega$ can be decomposed into $K$ clusters based on the image intensity $I:\Omega\to\mathbb{R}$, using the $K$-means clustering. Note that each cluster $\mathcal{K}_{i}$, may consist of multiple isolated components. We denote the isolated components by $D^{i}_{j}$. Hence, $\mathcal{K}_{i}=\bigcup_{j=1}^{N_{i}}D^{i}_{j}$, where $N_{i}$ indicates the number of isolated regions within cluster $i$. For each $\bm{x}_{\text{click}}\in\mathcal{C}_{*}$, we define $\delta(\bm{x}_{click})$ as follows:
$\delta(\bm{x}_{click})={D^{i}_{j}}\text{ for }(i,j)\text{ if }\bm{x}_{click}\in D^{i}_{j}.$
A local homogeneous region surrounding clicks can be obtained by taking the union, which is given by $\mathcal{H}=\bigcup_{\bm{x}_{click}\in\mathcal{C}_{*}}\delta(\bm{x}_{click})$. Then a click map $M_{\text{click}}$, which is a binary map capturing $\mathcal{H}$, can be defined as follows.
Definition 4.1 (Click map).
The click map $M_{\text{clicks}}$ for clicks $\mathcal{C}_{*}$ is defined as:
(5) $M_{\text{clicks}}(\bm{x})=\begin{cases}1,&\text{if }\bm{x}\in\mathcal{H},\\ 0,&\text{otherwise}.\end{cases}$
4.3 Fidelity term for interactive segmentation
The click map provides additional information to refine the segmentation prediction. In this subsection, we will analyze the properties of the fidelity term of the proposed interactive segmentation, which prompts us how to choose the suitable weight $r$.
4.3.1 Revisit on the fidelity term of (3)